There is a linear relationship between a woman's height and the length of her radius bone. It can be stated this way: Height increases by 3.9 inches for each 1-inch increase in the length of the radius. Suppose a 62-inch-tall woman has a 10-inch-long radius bone. Use this information to write a linear equation that relates height h to the length r of the radius. Write the equation in slope-intercept form.
Answer by TimothyLamb(4379)

slope = dy/dx
slope = (3.9)/(1)
slope = 3.9
---
h - 62 = 3.9(r - 10)
h = 3.9(r - 10) + 62
h = 3.9r - 39 + 62
h = 3.9r + 23
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---